A body vibrating with a simple harmonic motion has a frequency of 4hz and Amplitude of 0.15m. Calculate the maximum value of acceleration and velocity.

b) the acceleration and velocity at a point 0.09 from the equilibrium position of motion


1
Expert's answer
2021-01-12T12:12:34-0500

Solution:

"a_{max}\n\u200b\t\n =(2\u03c0f) \n^2\n A=(2(4)\u03c0)^2\n 0.15=95(\\tfrac{m}{s^2})"


"v_{max}\n\u200b\t\n =(2\u03c0f)A=(2(4)\u03c0)0.15=3.8(\\tfrac{m}{s})"


For the second part:


"a=(2(4)π) 
^2
 0.09=57(\tfrac{m}{s^2})"


"v=(2(4)π)0.09=2.3(\tfrac{m}{s})"